1/4

Christian Morgenstern, Christian Morgen- stern, Christian Morgenstern, Christian Morgen- stern, Christian Morgenstern,

The Bans- hee (An Approach) La lupfanto- mo Le loup-garou El He- chicero The Were- wolf

tradukita de Max Knight tradukita de Bertilo Wennergren tradukita de R. Platteau tradukita de Hans Eiseneck tradukita de Alexander Gross

One night, a banshee slunk away

from mate and child, and in the gloom

went to a village teacher’s tomb,

requesting him: “Inflect me, pray.”

Jen lupfantom’ de famili’ meznokte

iam kuris for al tombo de la profesor’. kaj petis pri konjugaci’.

Un loup-garou, une certaine nuit, Laissa sa moitié et puis se rendit Auprès de la tombe d’un maître d’école Et lui dit, “ sur moi dis tout, c’est ton rôle ! ”

Un hechicero, con esposa vino un día a la fosa de su maestro, y al profesor pidió: “¿Me cuentas? ¡Por favor!”

A Werewolf, troubled by his name, Left wife and brood one night and came

To a hidden graveyard to enlist The aid of a longdead philologist.

The village teacher climbed up straight

upon his grave stone with its plate

and to the appa-

rition said who meekly

knelt before the dead:

Do diktis tiu dokta ul’ starante sur la kructabul’ dum lup’ senmova kiel splinto atentis vortojn de l’ mortinto.

Un loup-garou, une certaine nuit, Laissa sa moitié et puis se rendit Auprès de la tombe d’un maître d’école Et lui dit, “ sur moi dis tout, c’est ton rôle ! ”

El profesor sin vacilar subió y comenzó a hablar. Desde allá de su letrero se dirigió al hechicero:

“Oh sage, wake up, please don’t berate me,” He howled sadly, “Just conjugate me.”

The seer arose a bit unsteady Yawned twice, wheezed once, and then was ready.

…  …  …  …  …

2/4

“The banSHEE, in the subject’s place;

the banHERS, the possessive case.

The banHER, next, is what they call

objective case– and that is all.”

De “lupfantomo”, simpla rekta,

tra “lupfintomo”, plej perfekta,

plu “lupfontomo”, iom rara,

ĝis “lupfuntomo”, ho, koŝmara.

“ - On peut à ton sujet puisque nous y sommes

Se demander encor bien des choses en somme.

Loup-garoù ? mais

avant Il faut se dire : loup-

garquand ?

“El hechiuno”

le explica, “el hechidos”

después indica,

“el hechitres”

dice al fin, “con eso

basta ¡chiquitín!”

“Well, ‘Werewolf’ is your plural past, While ‘Waswolf’ is singularly cast: There’s ‘Amwolf’ too, the present tense,

And ‘Iswolf,’ ‘Arewolf’ in this same sense.”

The banshee marveled at the cases

and writhed with pleasure, making faces,

but said: “You did not add, so far,

the plural to the singular!”

Ĉi participoj plaĉis multe

al la fantomo, kaj rezulte

ĝi petis: “Legu plu el PIV-o

pri participoj en pasivo.”

Loup-garcomment ? et

puis Bien évidemment

loup-garqui ? ” Toutes ces questions

semblèrent au loup bien absconses,

Car en fait ce qu’il voulait c’était les réponses.

El hechicero, conmovido

por lo que capta su oído,

al maestro ruega, encantado,

que le adjunte un quebrado.

“I know that–I’m no mental cripple– The future form and participle Are what I crave,” the beast replied.

The scholar paused– again he tried:

The teacher, though, admitted then

that this was not within his ken.

“While bans are frequent”, he advised,

“A she cannot be plurized.”

La viro ĉagreniĝis tamen:

“Jam fermis mi per fin’ kaj amen,

ĉar netransiraj bestoj feblas

kaj la pasiv’ neniam eblas.”

“ - Point de réponses, dit le puits de savoir,

Etre imaginaire, nul

ne peut te voir ! ” A ces mots l’animal

éclata tout en larmes.

“ - Oh!, même les baisers de ma douce moitié Pero el profe-

sor confiesa

que nada sabe con certeza:

“Hay muchos hechos. Hechiceros

únicamente hay enteros.”

“A ‘Will-be-wolf?’ It’s just too long: ‘Shall-be-wolf?’

‘Has-been-wolf?’ Utterly wrong! Such words are wounds beyond all suture– I’m sorry, but you have no future.”

…  …  …  …  …

3/4

The banshee, rising clammily,

wailed: “What about my family?”

Then, being not a learned creature,

said humbly “Thanks” and left the teacher.

Sed lupfantomo pri transiro ja spertas pli ol eĉ vampiro. La lup’ revenis larmokula al sia famili’ ulula.

Ne pourrons plus me faire me sentir tout entier.

Rester un loup-garou est un sort qui m’alarme. ” Mais comme le loup était bien poli Il remercia et puis re-

partit.

El brujo se volvió muy triste: “Mi cara mitad, ¿no la viste?” Ya que no era sabio con gratitud se despidió.

The Werewolf knew better–his sons still slept At home, and homewards now he crept,

Happy, humble, without apology For such folly of phi-

lology.

…  …  …  …  …

4/4

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Anglan de Max Knight (civila nomo: Max Kühnel, ∗1909-06-08 – †1993-08-31).

Arg-825-1660 (2013-02-13 22:36:06)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html . Pri la tradukinto Max Knight vidu: http: // www. jbeilharz. de/ morgenstern/ morgenstern_ poems. html kaj http: // www. 123people. de/ ext/ frm? ti= ersonensuche% 20telefonbuch&search_ term= m x% 20k% C3% BChnel&search_ country= DE&st= suc e% 20nach% 20personen&target_ url= aHR0cDovL d3dy5saXRlcmF0dXJoYXVzLmF0L2luZGV4LnBocD9pZD02ODc4 .

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en Esperanton de Bertilo Wennergren.

Arg-825-1653 (2013-02-14 23:29:47)

Vidu la retejon http: // bertilow. com/ literaturo/ lupfantomo. html .

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Francan de R. Platteau.

Arg-825-1655 (2013-02-13 21:56:38)

Tiu ĉi francigo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en hispana de Hans Eiseneck.

Arg-825-1656 (2013-02-13 18:46:00)

Tiu ĉi trsduko en la kastilian lingvon troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Anglan de Alexander Gross.

Arg-825-1657 (2013-02-13 18:39:45)

Tiu  ĉi  angli-
go  troviĝas  en

http: // bertilow. com/ literaturo/ lupfantomo. html .